A long elastic bungie cord with a Youngs modulus of 30MPa has initial length of 20m and cross-section of 55mm2 Taking care to convert the units correctly, determine the length of the cord when a person weighing 55 Kg is hanging in static equilibrium. Take g as 9.81N/Kg
Answer #1

Cross section area = 55 mm2 = 5.5 x 10-5 m2

Y = 30 MPa

Force on the cord = mg = 55 x 9.81 = 539.55 N

Young's modulus is: Y = stress/strain = FL/al

30 × 106-539.55(20) 5.5× 10-51

=> 6.54m .0477

so, the length of the cord when the person hangs on it is:

L' = 20 + 6.54 = 26.54 m.